# Chirally-protected state manipulation by tuning one-dimensional statistics

### v2

@@ Abstract
Chiral symmetry is broken by typical interactions in lattice models, but the statistical interactions embodied in the anyon-Hubbard model are an exception. It is an example of a correlated hopping model in which chiral symmetry protects a degenerate zero-energy subspace. Complementary to the traditional approach of anyon braiding in real space, we adiabatically evolve the statistical parameter in the anyon-Hubbard model and we find non-trivial Berry phases and holonomies in this chiral subspace. The corresponding states possess stationary checkerboard pattern in their $N$-particle densities which are preserved under adiabatic manipulation. We give an explicit protocol for how these chirally-protected zero energy states can be prepared, observed, validated, and controlled.
@@
In two dimensions, Abelian braid anyons with fractional exchange statistics arise from the topological analysis of two-body interactions [1, 2, 3, 4, 5, 6], and non-Abelian anyons of various forms have been proposed as the working material for robust topological quantum computing protocols [7, 8]. However, non-standard exchange statistics are not an exclusively two-dimensional phenomenon. Their key features have been proposed and investigated in one-dimensional systems since the beginning of the field [1, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], leading to experimental proposals [20, 21, 22, 23, 24, 25] and recent realizations in Raman-coupled Bose-Einstein condensates [26, 27]. This, for the first time, opens the possibility to tune the statistical angle of anyons in an experimentally accessible platform.
@@
The anyon-Hubbard model provides the platform for this exploration [15, 28, 29, 30]. It realizes exchange statistics with a statistical angle $\theta$ that interpolates between bosons $\theta=0$ and fermions $\theta=\pi$ are implemented on a one-dimensional lattice using Floquet-manipulated Rb atoms in quantum gas microscopes [31] by density-dependent Peierls phases  [32, 33, 34]. Such phases lead to intriguing effects, including statistically induced phase transitions [15, 28, 35], quasi-condensation at finite momenta [15, 28], emerging Friedel oscillations [29, 36, 30], as well as asymmetrical transport and expansion dynamics. [37, 38].
@@
Although the anyons realized by the anyon-Hubbard model are Abelian, their non-standard exchange statistics reveals topological structures in configuration space which we propose to be exploited for non-Abelian state manipulation. To motivate this, consider that the density-dependent Peierls phases which implement the statistical interaction can induce synthetic magnetic fluxes through plaquettes in configuration space [31, 19]. In the anyon-Hubbard model, these fluxes provide a phase $\exp(\pm i\theta)$ depending on the order in which the particles exchange; see Fig. 1(a). Many models in which underlying canonical particles experience correlated hopping processes share this feature, including [39, 32, 38, 33, 40, 41, 42]. Such correlated hopping processes can be engineered to break parity and time-reversal symmetry [43]. However, correlated hopping models with only nearest-neighbor hopping processes, such as the anyon-Hubbard model, preserve the chiral symmetry associated with bipartite lattices, a symmetry that is broken when non-statistical interactions are included [44, 45]. As a result, theorems about bipartite spatial lattices [46, 47, 48, 49] can be generalized from real space to configuration space in models with chirally symmetric interactions and particle number conservation.
@@
In this article, we show that one-dimensional lattice anyons with only statistical interactions host a degenerate zero energy subspace protected by chiral symmetry. Adiabatically tuning the statistical angle $\theta$ from 00 to $2\pi$ varies all energy levels in the spectrum except for the zero-energy modes, as shown in Fig. 1(b1-b2). In this space, the variation generates nontrivial holonomies [50, 51], i.e.,  unitary transformations similar to braiding non-Abelian two-dimensional anyons around each other [8]. This scheme can be used for chiral-protected non-Abelian state preparation. Building on theorems of Lieb and Sutherland [46, 47], we find that these results hold for the experimentally accessible anyon-Hubbard model and extend to chirally symmetric correlated hopping models of bosons with particle number conservation. In these models, chiral symmetry equips the configuration space with a stationary checkerboard pattern that can be experimentally revealed by the $N$-body density correlations and the spatially off-diagonal correlation functions of zero energy states and the quasi-momentum distribution. We furthermore show how these zero energy states can be prepared from typical initial states and manipulated by steering the statistical angle.
@@
Figure 1: Configuration space representation of the anyon-Hubbard model and its chirally-protected zero energy subspace $\mathcal{H}_{0}$. (a) Bosonic configuration space for $N=2$ particles on $L=6$ sites. Each dot represents a bosonic number state with doubly-occupied sites (black) and singly-occupied (not black) and with positive chirality (black, red hashed) or negative chirality (plain blue). The edges represent hopping processes with amplitude $-J$ (thin), $-\sqrt{2}{J}$ (thick, dotted), and $-\sqrt{2}Je^{i\theta}$ (arrow). Boundary plaquettes have a flux with a density-dependent Peierls phase $\theta$ through them. (b1) Energy spectrum as a function of the statistical angle $\theta$. The zero energy subspace has dimension $d_{0}=3$ for $N=2$ and $L=6$ (red line) without avoided crossings (inset (b2)). (b3) Schematic rotation of $\mathcal{H}_{0}$ by cyclic manipulation of $\theta$. (c) The minimal degeneracy $d_{0}$ of $\mathcal{H}_{0}$.
@@
Chiral symmetry and correlated hopping.— For a particle-number conserving bosonic system, a chiral symmetry corresponds to a unitary operator $\hat{S}$ that anticommutes with the Hamiltonian $[\hat{S},\hat{H}]_{+}=-\hat{H}$ [52, 53, 54]. Such an operator is an involution $\hat{S}^{2}=1$ that partitions the finite-dimensional Hilbert space into chiral eigenspaces $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ with eigenvalues $\chi=\pm 1$ and dimensions $\dim\mathcal{H}_{\pm}=d_{\pm}$. $\hat{H}$ is anti-block diagonal in the chiral basis and has a symmetric spectrum. Therefore, the zero energy eigenspace $\mathcal{H}_{0}$ has dimension $d_{0}\geq|d_{+}-d_{-}|$ [46, 47, 48, 49, 55]. This minimal degeneracy of the zero energy subspace $d_{0}$ does not depend on the details of the Hamiltonian and is protected against perturbations that preserves chiral symmetry.
@@
where $\hat{n}_{k}=\hat{b}^{\dagger}_{k}\hat{b}_{k}$ is the number operator on site $k$. The operator $\hat{S}$ transforms bosonic operators as $\hat{S}\hat{b}_{k}\hat{S}=(-1)^{k}\hat{b}_{k}$ and acts like a local gauge transformation in configuration space that assigns opposite chirality to adjacent number states, i.e., number states that differ by a single hop have opposite chirality. See Fig. 1(a) for a depiction of $L=6$ and $N=2$, where the chiral operator partitions the $21$-site configuration space lattice into a checkerboard of sublattices with $d_{+}\!=\!12$ and $d_{-}\!=\!9$ chiral number states, respectively. Chiral symmetry together with particle-number conservation guarantees that this system has at least $d_{0}\!=\!|d_{+}-d_{-}|=3$ zero-energy states with positive chirality, although the specific subspace $\mathcal{H}_{0}\subset\mathcal{H}_{+}$ spanned by these three states depends on the Hamiltonian. Note that additional pairwise zero energy subspace degeneracies can only appear due to other non-Abelian symmetries or special accidental degeneracies. In particular, the non-interacting Bose-Hubbard model with $L=8$ and $N\geq 3$ has additional degeneracies at $E=0$ because of the trigonometric relation $\cos(\pi/9)+\cos(5\pi/9)+\cos(7\pi/9)=0$. These accidental, cyclotomic degeneracies are akin to the Pythagorean degeneracies of the infinite square well. They occur in pairs that split upon variation from $\theta=0$ and are therefore not chirally-protected.
On the zero energy subspace, the chiral operator $\hat{S}$ is promoted to an actual symmetry, because $\mathcal{H}_{0}$ is a subspace of the majority chirality eigenspace $\mathcal{H}_{\pm}$ [46, 47, 49, 55]. This implies that all correlation functions vanish which are not invariant with respect to $\hat{S}$. For example, the operator $\hat{b}^{\dagger}_{j}\hat{b}_{j+1}$ has odd chirality and therefore has a vanishing expectation value for every state in $\mathcal{H}_{0}$. Remarkably, we find for $N<L$ an analogous pattern in the $N$-particle densities. States in the zero energy subspace have support exclusively on the majority sublattice [46, 47, 56], forming an higher-dimensional generalization of a checkerboard pattern. The scenario of spontaneous symmetry breaking in the thermodynamic limit is ruled out by Elitzur’s theorem [57, 58, 59].
For general $N$ and $L$, the dimension of the chirally-protected zero energy subspace can be derived using combinatorics [49, 60, 61] and is depicted in Fig. 1(c):
@@
Model.— The statistical interactions of the anyon-Hubbard model are chirally symmetric, in contrast to ordinary Hubbard-type interactions. They embody momentum-dependent interactions that are periodic in the statistical angle $\theta$. The model is defined in terms of anyonic operators $\hat{a}_{j}$ with deformed commutation relations that obey fractional exchange statistics
$\hat{a}_{j}\hat{a}_{k}^{\dagger}-e^{-i\theta\mathrm{sgn}(j-k)}\hat{a}^{\dagger}_{k}\hat{a}_{j}$ $=$ $\delta_{jk}$
@@
This results in a correlated hopping process mediated by a density-dependent Peierls phase.
Properties of the zero energy subspace.— We investigate the properties of the zero energy subspace by exact diagonalization and find the dimension of the zero energy subspace $d_{0}$ typically assumes its minimal value according to Eq. (2) as expected when the only symmetries are Abelian. Additional accidental degeneracies appear for special $N$ and $L$ at $\theta=0$ that are not chirally-protected and unsuitable for adiabatic manipulation.
We first construct a convenient basis for the zero energy subspace at $\theta=0$ (non-interacting bosons) and use this as a experimentally-verifiable starting point for state manipulation. The single-particle operators
@@
have zero energy and chirality $\chi_{p}=+1$. There are precisely $d_{0}$ ways to distribute $N$ indistinguishable bosons among these chiral pair states $|p(\mu)\rangle$ and chiral single-particle states $|s\rangle$, and from these a standard basis for the zero energy subspace can be built.
Figure 2: Holonomies for adiabatic evolutions from $\theta=0$ to $2\pi$. (a) The holonomy matrix $\mathbf{G}$ in pair basis $|p(\mu)\rangle$ (Eq. (7)) reveals non-trivial adiabatic holonomies ( here $N=2$, $L=6$, and $N_{\theta}=10^{4}$ in Eq. (8), error $\sim\!10^{-3}$, see [62]). (b) The smallest integer $m_{\mathrm{diag}}$ such that $G^{m_{\mathrm{diag}}}$ is diagonal (within the error bounds) quantifies how nontrivial $\hat{G}$ is. $m_{\mathrm{diag}}$ increases strongly with system size. The two-body density of an initial zero energy state (c), here $|p(1)\rangle$, alters significantly during the adiabatic evolution (d).
Adiabatic manipulation of the zero energy subspace.— The periodic dependency in the statistical angle $\theta$ invites the question how the zero energy subspace is affected under cyclic adiabatic variation of $\theta$. As the statistical angle is tuned, the zero energy subspace $\mathcal{H}_{0}$ moves and rotates within the larger subspace $\mathcal{H}_{\pm}$ with majority chirality $\chi=\pm 1$. Similar to the Aharonov-Bohm effect on a ring, the zero energy subspace accumulates topological Berry phases and nontrivial holonomies. These act as unitary transformations on the above-given standard basis vectors when tuning $\theta$ from 00 to $2\pi$, as schematically depicted in Fig. 1(b3).
We propagate the zero energy subspace from $\theta=0$ to $\theta=2\pi$ using Kato’s adiabatic evolution [50, 63]
@@
where $\hat{P}_{j}=\sum_{\mu=1}^{d_{0}}|\psi_{\mu}(2\pi j/N_{\theta})\rangle\langle\psi_{\mu}(2\pi j/N_{\theta})|$ is the projector onto the zero energy subspace at the statistical angle $2\pi j/N_{\theta}$ and $|\psi_{\mu}\rangle$ is an orthonormal basis for the chiral zero energy subspace; see [62] for numerical details. The unitary matrix $\mathbf{G}_{\mu\mu^{\prime}}=\langle\psi_{\mu^{\prime}}|\hat{G}|\psi_{\mu}\rangle$ then embodies the adiabatic holonomy for a complete loop.
@@
For $L=2$ the zero energy subspace is non-degenerate and we calculate the (Abelian) Berry phase $\mathbf{G}=\pm 1$ exactly and find the nontrivial value of $-1$ for consecutive odd-integer coefficients in the Gauss sum $\sum_{k=1}^{N/2}k$, see [62]. We investigate the non-Abelian holonomies of larger systems by exact diagonalization [62, 64, 65]. As an example, in Fig. 2(a), we present the holonomy matrix $\mathbf{G}$ in the chiral pair basis $|p(\mu)\rangle$ at $\theta=0$ for a system with $N=2$ and $L=6$ and $d_{0}=3$. The adiabatic evolution causes a significant non-trivial rotation of the zero energy subspace during the adiabatic evolution, which we indicate by ${\rho}_{ij}^{\nu}\left(\theta\right)=\langle\nu(\theta)|\hat{b}_{i}^{\dagger}\hat{b}_{j}^{\dagger}\hat{b}_{j}\hat{b}_{i}|\nu(\theta)\rangle$, an observable accessible in the corresponding experiments [31]. For example, we can take the initial two-body density $\rho_{ij}^{\nu}(\theta=0)$ of the basis state $|p(1)\rangle$ [Fig. 2(c)] and compare with ${\rho}_{ij}^{\nu}(\theta=2\pi)$, i.e., the two-body density corresponding to the adiabatically propagated state $\hat{G}|p(1)\rangle$. This yields a significant difference in the densities, see Fig. 2(d). Interestingly, for variations of parameters other than $\theta$ that preserve chiral symmetry, such as local variations in hopping strength $J$, we find that the connection is flat, meaning that $\hat{G}$ is not altered. Therefore, the results for the holonomy matrix $\mathbf{G}$ are robust against fluctuations in the adiabatic manipulation.
@@
Tuning the statistical angle therefore implements a topologically protected operation on the zero energy subspace, a concept which has been extensively explored in the context of Majorana modes and other non-Abelian anyons [8]. To quantify the nontriviality of $\mathbf{G}$, we determine the smallest integer $m_{\mathrm{diag}}$ such that $\mathbf{G}^{m_{\mathrm{diag}}}$ is diagonal, i.e., trivial in the context of state manipulation. Also, at least $m_{\mathrm{diag}}$ states in the computational space can be prepared from an initial state by repeated application of this operation. For instance, the braiding of Majorana modes becomes diagonal for $m_{\text{Maj.}}=2$ and trivial with $G_{\text{Maj.}}^{4}=1$ [8]. Within our error bounds, we find $m_{\text{diag}}>4$ for various particle numbers and system sizes, particularly for large ones, see Fig. 2(b) and [62].
@@
This idea can be elevated to non-Abelian holonomies if we implement the site-dependent statistical parameter $\theta_{j}$. This leads to a generalized anyonic exchange algebra $a^{\dagger}_{j}a_{k}=\hat{a}_{i}\hat{a}_{j}^{\dagger}-e^{i\theta_{i,j}}\hat{a}_{j}^{\dagger}\hat{a}_{i}=\delta_{i,j}$ with $\theta_{i,j}=\theta_{j}$ for $i>j$, $\theta_{i,j}=-\theta_{i}$ for $i<j$, and $\theta_{i,i}=0$. For different exchange phases $\theta_{L}$ in the left and $\theta_{R}$ in the right part of the system [62, 66], we generally find $\left[\mathbf{G}_{L},\mathbf{G}_{R}\right]\neq 0$, meaning the fluctuation-protected state manipulation depends on which statistical angle is altered first [62].
@@
Figure 3: Steering to the zero energy subspace, starting with $N=2$ bosons at the center of $L=6$ sites. (a) The statistical phase $\theta(t)$ is dynamically varied along $60$ optimized interpolation points (gray) such that (b) the overlap $\mathcal{O}(t_{f})$ of the time-evolved state with the zero energy state $|p(1)\rangle$ (Eq. (7)) is maximal, reaching a final overlap of $1-10^{-4}$. The time evolution of (c) the chirality and (d) the static structure factor $C_{\pi}(t)$, which approaches its maximal value $N^{2}/L^{2}$ (dashed line) indicating the emergence of a checkerboard pattern, the signature of the zero energy subspace.
@@
Steering to the zero energy subspace.— The basis states for $\theta=0$ can be prepared by temporal variation of $\theta=\theta(t)$ [31]. As a simple, accessible initial state, we consider two bosonic particles at the central sites of a lattice with $L=6$ sites, i.e., $|\Psi(t=0)\rangle=|001100\rangle$. In order to find a path for $\theta(t)$ which maximizes the overlap of the propagated wave function at time $t_{f}$ with a chiral pair target state, $\mathcal{O}(t_{f})=|\langle p(\mu)|\Psi(t_{f})\rangle|^{2}$, we represent $\theta(t)$ with $M$ interpolation points (for details see [62]). These $M$ points serve as an input for a gradient based optimization algorithm which updates the path $\theta(t)$ until $\mathcal{O}(t_{f})$ becomes maximal [67, 68, 69].

### v3

@@ Abstract
Chiral symmetry is broken by typical interactions in lattice models, but the statistical interactions embodied in the anyon-Hubbard model are an exception. This is an example for a correlated hopping model where chiral symmetry protects a degenerate zero-energy subspace. Complementary to the traditional approach of anyon braiding in real space, we adiabatically evolve the statistical parameter and find non-trivial Berry phases and holonomies in this chiral subspace. The corresponding states possess stationary checkerboard patterns in their $N$-particle densities which are preserved under adiabatic manipulation. We give an explicit protocol for how these chirally-protected zero-energy states can be prepared, observed, validated, and controlled.
@@
In two dimensions, Abelian braid anyons with fractional exchange statistics arise from the topological analysis of configuration spaces [1, 2, 3, 4, 5, 6], and non-Abelian anyons of various forms have been proposed as the working material for robust topological quantum computing protocols [7, 8]. However, non-standard exchange statistics are not an exclusively two-dimensional phenomenon. Their key features have been proposed and investigated in one-dimensional systems since the beginning of the field [1, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], leading to experimental proposals [20, 21, 22, 23, 24, 25] and recent realizations in Raman-coupled Bose-Einstein condensates [26, 27]. This, for the first time, opens the possibility to tune the statistical angle of anyons in an experimentally accessible platform.
@@
The anyon-Hubbard model offers the possibility for this exploration [15, 28, 29, 30]. Exchange statistics described by a statistical angle $\theta$ that interpolates between bosons $\theta=0$ and fermions $\theta=\pi$ are implemented on a one-dimensional lattice using Floquet-manipulated Rb atoms in quantum gas microscopes [31] by density-dependent Peierls phases  [32, 33, 34]. Such phases lead to intriguing effects, including statistically induced phase transitions [15, 28, 35], quasi-condensation at finite momenta [15, 28], emerging Friedel oscillations [29, 36, 30], as well as asymmetrical transport and expansion dynamics. [37, 38].
@@
Although the anyons realized by the anyon-Hubbard model are Abelian, their non-standard exchange statistics reveals topological structures in configuration space which we propose to exploit for non-Abelian state manipulation. To motivate this, consider that the density-dependent Peierls phases which implement the statistical interaction can induce synthetic magnetic fluxes through plaquettes in configuration space [31, 19]. In the anyon-Hubbard model, these fluxes provide a phase $\exp(\pm i\theta)$ depending on the order in which the particles exchange; see Fig. 1(a). Many models in which underlying canonical particles experience correlated hopping processes share this feature, including [39, 32, 38, 33, 40, 41, 42]. Such correlated hopping processes can be engineered to break parity and time-reversal symmetry [43]. However, correlated hopping models with nearest-neighbor hopping processes, such as the anyon-Hubbard model without on-site, Hubbard interactions, preserve the chiral symmetry associated with bipartite lattices [44, 45]. As a result, theorems about bipartite spatial lattices [46, 47, 48, 49] can be generalized from real space to configuration space in models with chirally symmetric interactions and particle number conservation.
@@
In this article, we show that one-dimensional lattice anyons with only statistical interactions host a degenerate zero-energy subspace protected by chiral symmetry. Adiabatically tuning the statistical angle $\theta$ from 00 to $2\pi$ varies all energy levels in the spectrum except for the zero-energy modes, as shown in Fig. 1(b1-b2). In this space, the variation generates nontrivial holonomies [50, 51], i.e., unitary transformations similar to braiding non-Abelian two-dimensional anyons around each other [8]. This scheme can be used for chirally-protected non-Abelian state preparation. Building on theorems of Lieb and Sutherland [46, 47], we find that these results hold for the anyon-Hubbard model in the experimentally accessible limit of only statistical interactions [31] and extend to chirally-symmetric correlated hopping models of bosons with particle number conservation. In these models, chiral symmetry equips the configuration space with a stationary checkerboard pattern that can be experimentally revealed by the $N$-body density correlations and the spatially off-diagonal correlation functions of zero-energy states and the quasi-momentum distribution. We furthermore show how these zero-energy states can be prepared from typical initial states and manipulated by steering the statistical angle.
@@
Figure 1: Configuration space representation of the anyon-lattice model and its chirally-protected zero-energy subspace $\mathcal{H}_{0}$. (a) Bosonic configuration space for $N=2$ particles on $L=6$ sites. Each dot represents a bosonic number state with doubly-occupied sites (black) and singly-occupied (not black) and with positive chirality (black, red hashed) or negative chirality (plain blue). The edges represent hopping processes with amplitude $-J$ (thin), $-\sqrt{2}{J}$ (thick, dotted), and $-\sqrt{2}Je^{i\theta}$ (arrow). Boundary plaquettes have a flux with a density-dependent Peierls phase $\theta$ through them. (b1) Energy spectrum as a function of the statistical angle $\theta$. The zero-energy subspace has dimension $d_{0}=3$ for $N=2$ and $L=6$ (red line) without avoided crossings (inset (b2)). (b3) Schematic rotation of $\mathcal{H}_{0}$ by cyclic manipulation of $\theta$. (c) The minimal degeneracy $d_{0}$ of $\mathcal{H}_{0}$.
@@
Chiral symmetry and correlated hopping.— For a particle-number conserving bosonic system, a chiral symmetry corresponds to a unitary operator $\hat{S}$ that anticommutes with the Hamiltonian $[\hat{S},\hat{H}]_{+}=0$ [52, 53, 54]. Such an operator is an involution $\hat{S}^{2}=1$ that partitions the finite-dimensional Hilbert space into chiral eigenspaces $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ with eigenvalues $\chi=\pm 1$ and dimensions $\dim\mathcal{H}_{\pm}=d_{\pm}$. $\hat{H}$ is anti-block diagonal in the chiral basis and has a symmetric spectrum. Therefore, the zero-energy eigenspace $\mathcal{H}_{0}$ has dimension $d_{0}\geq|d_{+}-d_{-}|$ [46, 47, 48, 49, 55, 56]. This minimal degeneracy of the zero-energy subspace $d_{0}$ does not depend on the details of the Hamiltonian and is protected against perturbations that preserves chiral symmetry.
@@
where $\hat{n}_{k}=\hat{b}^{\dagger}_{k}\hat{b}_{k}$ is the number operator on site $k$. The operator $\hat{S}$ transforms bosonic operators as $\hat{S}\hat{b}_{k}\hat{S}=(-1)^{k}\hat{b}_{k}$ and acts like a local gauge transformation in configuration space that assigns opposite chirality to adjacent number states, i.e., number states that differ by a single hop have opposite chirality. See Fig. 1(a) for a depiction of $L=6$ and $N=2$, where the chiral operator partitions the $21$-site configuration space lattice into a checkerboard of sublattices with $d_{+}\!=\!12$ and $d_{-}\!=\!9$ chiral number states, respectively. Chiral symmetry together with particle-number conservation guarantees that this system has at least $d_{0}\!=\!|d_{+}-d_{-}|=3$ zero-energy states with positive chirality, although the specific subspace $\mathcal{H}_{0}\subset\mathcal{H}_{+}$ spanned by these three states depends on the Hamiltonian. Note that additional pairwise zero-energy degeneracies can appear due to other non-Abelian symmetries or special accidental degeneracies. In particular, the non-interacting Bose-Hubbard model with $L=8$ and $N\geq 3$ has additional degeneracies at $E=0$ because of the trigonometric relation $\cos(\pi/9)+\cos(5\pi/9)+\cos(7\pi/9)=0$. These accidental, cyclotomic degeneracies are akin to the Pythagorean degeneracies of the infinite square well. They occur in pairs that split upon variation from $\theta=0$ and are therefore not chirally-protected.
On the zero-energy subspace, the chiral operator $\hat{S}$ is promoted to an actual symmetry, because $\mathcal{H}_{0}$ is a subspace of the majority chirality eigenspace $\mathcal{H}_{\pm}$ [46, 47, 49, 55, 56]. This implies that all correlation functions vanish which are not invariant with respect to $\hat{S}$. For example, the operator $\hat{b}^{\dagger}_{j}\hat{b}_{j+1}$ has odd chirality and therefore has a vanishing expectation value for every state in $\mathcal{H}_{0}$ leading to the emergence of a checkerboard pattern that is stationary in time. Remarkably, we find for $N<L$ an analogous pattern in the $N$-particle densities. States in the zero-energy subspace have support exclusively on the majority sublattice [46, 47, 57], forming an higher-dimensional generalization of a checkerboard pattern. The scenario of spontaneous symmetry breaking in the thermodynamic limit is ruled out by Elitzur’s theorem [58, 59, 60].
For general $N$ and $L$, the dimension of the chirally-protected zero-energy subspace can be derived using combinatorics [49, 61, 62] and is depicted in Fig. 1(c):
@@
Model.— The statistical interactions of the anyon-lattice model are chirally symmetric, in contrast to ordinary Hubbard-type interactions. They embody momentum-dependent interactions that are periodic in the statistical angle $\theta$. The model is defined in terms of anyonic operators $\hat{a}_{j}$ with deformed commutation relations that obey fractional exchange statistics
$\hat{a}_{j}\hat{a}_{k}^{\dagger}-e^{-i\theta\mathrm{sgn}(j-k)}\hat{a}^{\dagger}_{k}\hat{a}_{j}$ $=$ $\delta_{jk},$
@@
This results in a correlated hopping process mediated by a density-dependent Peierls phase. The $U=0$ instantiation of the anyon-Hubbard model (5) has been successfully implemented in a Floquet-driven quantum gas microscope [31].
Properties of the zero-energy subspace.— We investigate the properties of the zero-energy subspace by exact diagonalization and find the dimension of the zero-energy subspace $d_{0}$ typically assumes its minimal value according to Eq. (2) as expected when the only symmetries are Abelian. Additional accidental degeneracies appear for special $N$ and $L$ at $\theta=0$ that are not chirally-protected and unsuitable for adiabatic manipulation.
We first construct a convenient basis for the zero-energy subspace at $\theta=0$ (non-interacting bosons) and use this as a experimentally-verifiable starting point for state manipulation. The single-particle operators
@@
have zero energy and chirality $\chi_{p}=+1$. There are precisely $d_{0}$ ways to distribute $N$ indistinguishable bosons among these chiral pair states $|p(\mu)\rangle$ and chiral single-particle states $|s\rangle$, and from these a standard basis for the zero-energy subspace can be built. As expected, these states exhibit a stationary checkerboard pattern in configuration space.
Figure 2: Holonomies for adiabatic evolutions from $\theta=0$ to $2\pi$. (a) The holonomy matrix $\mathbf{G}$ in pair basis $|p(\mu)\rangle$ (Eq. (7)) reveals non-trivial adiabatic holonomies ( here $N=2$, $L=6$, and $N_{\theta}=10^{4}$ in Eq. (8), error $\sim\!10^{-3}$, see [63]). (b) The smallest integer $m_{\mathrm{diag}}$ such that $G^{m_{\mathrm{diag}}}$ is diagonal (within the error bounds) quantifies how nontrivial $\hat{G}$ is. $m_{\mathrm{diag}}$ increases strongly with system size. The two-body density of an initial zero-energy state (c), here $|p(1)\rangle$, alters significantly during the adiabatic evolution (d).
Adiabatic manipulation of the zero-energy subspace.— The periodic dependency in the statistical angle $\theta$ invites the question how the zero-energy subspace is affected under cyclic adiabatic variation of $\theta$. As the statistical angle is tuned, the zero-energy subspace $\mathcal{H}_{0}$ moves and rotates within the larger subspace $\mathcal{H}_{\pm}$ with majority chirality $\chi=\pm 1$. Similar to the Aharonov-Bohm effect on a ring, the zero-energy subspace accumulates topological Berry phases and nontrivial holonomies. These act as unitary transformations on the above-given standard basis vectors when tuning $\theta$ from 00 to $2\pi$, as schematically depicted in Fig. 1(b3).
We propagate the zero-energy subspace from $\theta=0$ to $\theta=2\pi$ using Kato’s adiabatic evolution [50, 64]
@@
where $\hat{P}_{j}=\sum_{\mu=1}^{d_{0}}|\psi_{\mu}(2\pi j/N_{\theta})\rangle\langle\psi_{\mu}(2\pi j/N_{\theta})|$ is the projector onto the zero-energy subspace at the statistical angle $2\pi j/N_{\theta}$ and $|\psi_{\mu}\rangle$ is an orthonormal basis for the chiral zero-energy subspace; see [63] for numerical details. The unitary matrix $\mathbf{G}_{\mu\mu^{\prime}}=\langle\psi_{\mu^{\prime}}|\hat{G}|\psi_{\mu}\rangle$ then embodies the adiabatic holonomy for a complete loop.
@@
For $L=2$ and $N$ even where the zero-energy subspace is non-degenerate, we can exactly calculate the (Abelian) holonomy to $\mathbf{G}=\exp\left[i\pi N(N/2+1)/4\right]=\pm 1$, see [63]. More generally, we must investigate the non-Abelian holonomies of larger systems by exact diagonalization [63, 65, 66]. As an example, in Fig. 2(a), we present the holonomy matrix $\mathbf{G}$ in the chiral pair basis $|p(\mu)\rangle$ at $\theta=0$ for a system with $N=2$ and $L=6$ and $d_{0}=3$. The adiabatic evolution causes a significant non-trivial rotation of the zero-energy subspace during the adiabatic evolution, which we indicate by ${\rho}_{ij}^{\nu}\left(\theta\right)=\langle\nu(\theta)|\hat{b}_{i}^{\dagger}\hat{b}_{j}^{\dagger}\hat{b}_{j}\hat{b}_{i}|\nu(\theta)\rangle$, an observable accessible in the corresponding experiments [31]. Additionally, note that an adiabatic scheme has been realized in a similar setup [67] and the experiment reported in [31] demonstrated a high control over the on-site interaction strength. We have checked that the holonomy matrix in Fig. 2(a) is robust against small modulations of possible on-site interactions (see [63]). For example, we can take the initial two-body density $\rho_{ij}^{\nu}(\theta=0)$ of the basis state $|p(1)\rangle$ [Fig. 2(c)] and compare with ${\rho}_{ij}^{\nu}(\theta=2\pi)$, i.e., the two-body density corresponding to the adiabatically propagated state $\hat{G}|p(1)\rangle$. This yields a significant difference in the densities, see Fig. 2(d). Interestingly, for variations of parameters other than $\theta$ that preserve chiral symmetry, such as local variations in hopping strength $J$, we find that the connection is flat, meaning that $\hat{G}$ is not altered. Consequently, the results for the holonomy matrix $\mathbf{G}$ are robust against fluctuations in the adiabatic manipulation. Moreover, the zero-energy subspace gets larger as $N$ and $L$ increase, allowing for a richer structure of non-Abelian holonomies. However, the energy gap between the zero-energy states and the adjacent energy levels also narrows, which increases the adiabatic time scale and enhances incoherence by diabatic noise.
@@
Tuning the statistical angle therefore implements a topologically protected operation on the zero-energy subspace, a concept which has been extensively explored in the context of Majorana modes and other non-Abelian anyons [8]. To quantify the nontriviality of $\mathbf{G}$, we determine the smallest integer $m_{\mathrm{diag}}$ such that $\mathbf{G}^{m_{\mathrm{diag}}}$ is diagonal, i.e., trivial in the context of state manipulation. Also, at least $m_{\mathrm{diag}}$ states in the computational space can be prepared from an initial state by repeated application of this operation [68]. For instance, the braiding of Majorana modes becomes diagonal for $m_{\text{Maj.}}=2$ and trivial with $G_{\text{Maj.}}^{4}=1$ [8]. Within our error bounds, we find $m_{\text{diag}}>4$ for various particle numbers and system sizes, particularly for large ones, see Fig. 2(b) and [63].
@@
This idea can be elevated to non-Abelian holonomies if we implement the site-dependent statistical parameter $\theta_{j}$. This leads to a generalized anyonic exchange algebra $a^{\dagger}_{j}a_{k}=\hat{a}_{i}\hat{a}_{j}^{\dagger}-e^{i\theta_{i,j}}\hat{a}_{j}^{\dagger}\hat{a}_{i}=\delta_{i,j}$ with $\theta_{i,j}=\theta_{j}$ for $i>j$, $\theta_{i,j}=-\theta_{i}$ for $i<j$, and $\theta_{i,i}=0$. For different exchange phases $\theta_{L}$ in the left and $\theta_{R}$ in the right part of the system [63, 69], we generally find non-Abelian holonomies with $\left[\mathbf{G}_{L},\mathbf{G}_{R}\right]\neq 0$, meaning the fluctuation-protected state manipulation depends on which statistical angle is altered first [63].
@@
Figure 3: Steering to the zero-energy subspace, starting with $N=2$ bosons at the center of $L=6$ sites. (a) The statistical phase $\theta(t)$ is dynamically varied along $60$ optimized interpolation points (gray) such that (b) the overlap $\mathcal{O}(t_{f})$ of the time-evolved state with the zero-energy state $|p(1)\rangle$ (Eq. (7)) is maximal, reaching a final overlap of $1-10^{-4}$. The time evolution of (c) the chirality and (d) the static structure factor $C_{\pi}(t)$, which approaches its maximal value $N^{2}/L^{2}$ (dashed line) indicating the emergence of a checkerboard pattern, the signature of the zero-energy subspace.
@@
Steering to the zero-energy subspace.— The basis states for $\theta=0$ can be prepared by temporal variation of $\theta=\theta(t)$ [31]. As a simple, accessible initial state, we consider two bosonic particles at the central sites of a lattice with $L=6$ sites, i.e., $|\Psi(t=0)\rangle=|001100\rangle$. In order to find a path for $\theta(t)$ which maximizes the overlap of the propagated wave function at time $t_{f}$ with a chiral pair target state, $\mathcal{O}(t_{f})=|\langle p(\mu)|\Psi(t_{f})\rangle|^{2}$, we represent $\theta(t)$ with $M$ interpolation points (for details see [63]). These $M$ points serve as an input for a gradient based optimization algorithm which updates the path $\theta(t)$ until $\mathcal{O}(t_{f})$ becomes maximal [70, 71, 72].